# Page Curve and Entanglement Dynamics in an Interacting Fermionic Chain

### v2

@@ Appendix A Entanglement Hamiltonian
${\cal F}_{E}(\beta_{\text{fictitious}})=-\frac{1}{\beta_{\text{fictitious}}}{\cal Z}_{E}(\beta_{\text{fictitious}})$ (15)
@@
Here $\beta_{\text{fictitious}}=1/T_{\text{fictitious}}$ is the inverse of the fictitious temperature. Comparing against the Rényi entropy $S_{n}$ in Eq. (5), we find a proportionality between the Rényi entropy and the thermodynamic free energy (thereby connecting the two)
$S_{n}=\frac{n}{n-1}{\cal F}_{E}(n)$ (17)
which implies that $\beta_{\text{fictitious}}=n$ and therefore concludes our discussion that the fictitious temperature of the entanglement Hamiltonian is given by $1/n$.
@@ Appendix B Additional plots for different interacting strengths
Next we analyze the entanglement Hamiltonian for these two interacting strengths where we again find an underlying quantum phase transition. We plot the eigenvalues and the Schmidt coefficients (using which the eigenvalues of the entanglement Hamiltonian are calculated) for $V=0.4$ in Fig. 15 against the decayed fraction of particles into the environment, as well as against normalized time in Fig. 16.
@@
Figure 17: We plot the quantum phase transition that results in non-analyticity in min-entropy for $V=1.2$. The rest of the parameters as well as the conclusions are the same as stated in the captions of Figs. 6 and 15.
@@
Table 2: Linear regression for the first non-analytic points observed in min-entropy (see Figs. 14) observed across different system sizes ($M=3$ to $M=7$ for a fixed total number of lattice sites $L=50$) that corresponds to the quantum phase transition in the entanglement Hamiltonian (crossing shown in Figs. 16 and 18). In the regression, the $Y-$axis is $t_{c}/M$ where $t_{c}$ is the critical time at which the non-analyticity develops in min-entropy. The $X-$axis is $1/M$, therefore the $Y-$intercept is the extrapolation of the critical time in the thermodynamic limit. For intermediate to strong interactions, the thermodynamic value for critical time vanishes, and we provide an analysis in the text based on Fig. 9. We also note an acute sensitivity on the tunneling strength $g$ where the thermodynamic value vanishes at intermediate interaction strength for larger values of $g$.
@@ Appendix C Strongly Interacting Case
As discussed in the analysis in Section V in the main text, in the strongly interacting case where $V/t_{h}\gg 1$ (where $t_{h}\equiv t_{s}=t_{e}$ which has been set to $1.0$ in this work), a maximum of one particle can leak into the environment. Accordingly, we were able to write down an ansatz for the number of particles in the system (Eq. (9)) where only the largest ($1-\lambda(t)$) and the second-largest ($\lambda(t)$) Schmidt coefficients play a role such that $\lambda(t=0)=0$. We show the results for $V=3.0$ and $V=5.0$ in Fig. 20 that indeed the system gets stuck despite showing short time fluctuations and there is no further leakage of particle, at least for the timescale accessible to us. We further plot in Fig. 21 the four largest Schmidt coefficients where we find (as claimed in the main text) that only the largest and the second-largest Schmidt coefficients play a dominating role in determining the temporal dynamics of the system as the interaction strength is further increased.
Figure 20: Results for strongly interacting cases for (a) V=3.0 and (b) V=5.0 where other parameter values are mentioned in the plot. As explained in the main text in Section V, the system gets stuck despite showing short term fluctuations.
Figure 21: The corresponding Schmidt coefficients for the graphs shown in Fig. 20 for strongly interacting cases for (a) V=3.0 and (b) V=5.0 where other parameter values are mentioned in the plot. As explained in the main text in Section V, only the largest and the second-largest Schmidt coefficients play a role which becomes evident in these plots as well. This allows us to write the ansatz for the particle number in the system, as in Eq. (9).

### v3

@@ Appendix A Entanglement Hamiltonian
${\cal F}_{E}(\beta_{\text{fictitious}})=-\frac{1}{\beta_{\text{fictitious}}}\ln{\cal Z}_{E}(\beta_{\text{fictitious}})$ (19)
@@
Here $\beta_{\text{fictitious}}=1/T_{\text{fictitious}}$ is the inverse of the fictitious temperature. Comparing against the Rényi entropy $S_{n}$ in Eq. (5), we find a proportionality between the Rényi entropies and the thermodynamic free energy of the entanglement Hamiltonian (thereby connecting the two) when the inverse fictitious temperature $\beta_{\text{fictitious}}=n$
$S_{n}=\frac{n}{n-1}{\cal F}_{E}(n).$ (21)
This concludes our discussion that the fictitious temperature of the entanglement Hamiltonian is given by $1/n$.
@@ Appendix B Additional plots for different interacting strengths
Next we analyze the entanglement Hamiltonian for these two interacting strengths where we again find an underlying quantum phase transition. We plot the eigenvalues and the Schmidt coefficients (using which the eigenvalues of the entanglement Hamiltonian are calculated) for V= 0.4 (V= 1.2) in Fig. 17 (19) against the decayed fraction of particles into the environment, as well as against normalized time in Figs. 18 and 20.
@@
Figure 19: We plot the quantum phase transition that results in non-analyticity in min-entropy for $V=1.2$. The rest of the parameters as well as the conclusions are the same as stated in the captions of Figs. 6 and 17.
@@
Table 4: Linear regression for the first non-analytic points observed in min-entropy (see Figs. 16) observed across different system sizes ($M=3$ to $M=7$ for a fixed total number of lattice sites $L=50$) that corresponds to the quantum phase transition in the entanglement Hamiltonian (crossing shown in Figs. 18 and 20). In the regression, the $Y-$axis is $t_{c}/M$ where $t_{c}$ is the critical time at which the non-analyticity develops in min-entropy. The $X-$axis is $1/M$, therefore the $Y-$intercept is the extrapolation of the critical time in the thermodynamic limit. For intermediate to strong interactions, the thermodynamic value for critical time vanishes, and we provide an analysis in the text based on Fig. 11. We also note an acute sensitivity on the tunneling strength $g$ where the thermodynamic value vanishes at intermediate interaction strength for larger values of $g$.
@@ Appendix C Strongly Interacting Case
As discussed in the analysis in Section VI in the main text, in the strongly interacting case where $V/t_{h}\gg 1$ (where $t_{h}\equiv t_{s}=t_{e}$ which has been set to $1.0$ in this work), a maximum of one particle can leak into the environment. Accordingly, we were able to write down an ansatz for the number of particles in the system (Eq. (13)) where only the largest ($1-\lambda(t)$) and the second-largest ($\lambda(t)$) Schmidt coefficients play a role such that $\lambda(t=0)=0$. We show the results for $V=3.0$ and $V=5.0$ in Fig. 22 that indeed the system gets stuck despite showing short time fluctuations and there is no further leakage of particle, at least for the timescale accessible to us. We further plot in Fig. 23 the four largest Schmidt coefficients where we find (as claimed in the main text) that only the largest and the second-largest Schmidt coefficients play a dominating role in determining the temporal dynamics of the system as the interaction strength is further increased. As mentioned above Eq. (13), strong repulsive force not necessarily lead to particles dispersing, but the dynamics can completely freeze in the presence of strong interactions. This is what we observe here.
Figure 22: Results for strongly interacting cases for (a) V=3.0 and (b) V=5.0 where other parameter values are mentioned in the plot. As explained in the main text in Section VI, the system gets stuck despite showing short term fluctuations.
Figure 23: The corresponding Schmidt coefficients for the graphs shown in Fig. 22 for strongly interacting cases for (a) V=3.0 and (b) V=5.0 where other parameter values are mentioned in the plot. As explained in the main text in Section VI, only the largest and the second-largest Schmidt coefficients play a role which becomes evident in these plots as well. This allows us to write the ansatz for the particle number in the system, as in Eq. (13).
Appendix D Further Plots for Finite Size Scaling and Critical Exponent
We present here the plots for $V=0.0$, $V=0.4$ and $V=0.8$ in Figs. 24, 25 and 26. The descriptions of the top and the bottom rows in each of these plots are explained in the caption of Fig. 8.
Figure 24: Parameters: $V=0.8$, $g=0.5$, $t_{s}=t_{e}=1.0$, $L=50$ and $M=\{3,4,5,6,7\}$.
Figure 25: Parameters: $V=0.4$, $g=0.5$, $t_{s}=t_{e}=1.0$, $L=50$ and $M=\{3,4,5,6,7\}$.
Figure 26: Parameters: $V=0.0$, $g=0.5$, $t_{s}=t_{e}=1.0$, $L=50$ and $M=\{3,4,5,6,7\}$. This is also used for benchmarking as the free case is solvable analytically in the thermodynamic limit [27].